# Enhancing Maritime Trajectory Forecasting via H3 Index and Causal Language Modelling (CLM)

### v1

@@ 5.1 Metrics analysis
The violin plot in figure 3 shows a dense clustering of data points, meaning that the majority of the model’s predictions fall within lower Fréchet distance ranges. This indicates a high level of accuracy in predicting ship trajectories under a broad set of conditions. In practical terms, this suggests that the model is adept at capturing the essential patterns and dynamics of maritime movement in many scenarios, making it a reliable tool for planning and navigation in familiar waters or under standard operating conditions (see C and D).
The long tail extending into higher Fréchet distances highlights the limitations of the model in certain complex scenarios. These outliers represent instances where the predicted trajectory significantly deviates from the actual path taken by the ship. Such deviations could arise from unpredictable environmental factors (like sudden weather changes), complex maritime traffic conditions, or unique navigational challenges not fully encapsulated by the model’s parameters such as the trajectories of fishing boats, which are difficult to predict by taking only GNSS positions into account. These cases underscore the challenges of modelling the intricacies of real-world navigation, where unforeseen variables can impact a ship’s course.
Figure 3: Violin plot of the prediction distance without outliers, highlighting the mean (green star, indicated by the green arrow head), median, and density peak for a context length of 60 minutes and a prediction length of 90 minutes. See figures 11, 12 and 13 in E for the full analysis.
@@
The figure 4 highlights that the prediction error generally increases with the prediction distance, as shown by the comparison between (a) and (b). It is generally observed that the prediction errors are greater as the context length is reduced, and the prediction length is increased. On the other hand, we can see that, despite certain hallucinations which we will discuss later, the model seems to have generalised the nature of a trajectory, as shown in image (c) of figure 14 in F.
@@ 5.3 Hallucination
Figure 7: An example of the phenomena we have called "hallucinations". A red line represents the context used to generate the green trajectories. A blue line represents the ground truth. The red point is the barycentre of the polygon formed by the green points at the ends of the green trajectories.
Currently, there is no concrete explanation for the occurrence of these hallucinations. However, it is acknowledged that this phenomenon is exceedingly rare. Despite its infrequency, the ability to visually distinguish these erroneous outputs is crucial. Marine operation teams can easily identify these hallucinations, ensuring that such unrealistic predictions are not considered in practical navigation decisions.
@@ Appendix F Multiple trajectories grid
(a) Refer to caption Refer to caption Refer to caption
(b) Refer to caption Refer to caption Refer to caption
(c) Refer to caption Refer to caption Refer to caption

### v2

@@ 4.5 Qualitative and Quantitative Evaluation of Trajectory Predictions
Additionally, we conducted a quantitative comparison with TrAISformer, introduced in earlier sections as a state-of-the-art reference. To ensure consistency, we reproduced the TrAISformer experiment ourselves using the code provided by the original authors. However, we modified the evaluation metric to use the Fréchet distance, as discussed previously, and used TrAISformer’s evaluation dataset to measure the performance of our model, MixTRAJ, both with and without fine-tuning.
@@ 5.1 Metrics analysis
The violin plots in figure 5 depict the distribution of two metrics used to evaluate the performance of our model in predicting ship trajectories: the prediction distance and the Fréchet distance. Sub-figure (a) shows the distribution of the prediction distance (in metres), calculated as the straight-line distance between the last point of the context and the last point of the predicted trajectory. With a mean of approximately 29,790 metres and a density peak around 32,401 metres, this value represents the average distance covered in 90 minutes of prediction. This metric helps to contextualise the model’s output by quantifying the length of trajectory being predicted over time, as our model outputs tokens only.
In sub-figure (b), the Fréchet distance is used to compare the predicted trajectory with the actual one. Fréchet distance measures the similarity between the full predicted and actual paths, capturing the accuracy of the model over the entire trajectory rather than just at the endpoints. This distance is calculated using the closest ground-truth point to the centroid formed by the $N$ predictions in a given context. The mean Fréchet distance is 2,806 metres, with a median of 1,938 metres and a density peak at 1,061 metres. The proximity of the density peaks in both sub-figures suggests that the model performs consistently, with a typical prediction error of around 1 kilometre for an overall predicted distance of 32 kilometres. This demonstrates that the model effectively captures the global patterns of ship movement under most conditions (see appendix C and D).
The long tails in both plots highlight instances where the model’s predictions deviate more significantly from the actual trajectories. Such deviations could arise from unpredictable environmental factors (like sudden weather changes), complex maritime traffic conditions, or unique navigational challenges not fully encapsulated by the model’s parameters such as the trajectories of fishing boats, which are difficult to predict by taking only GNSS positions into account. These cases underscore the challenges of modelling the intricacies of real-world navigation, where unforeseen variables can impact a ship’s course.
Figure 5: Violin plot of the prediction distance. Without outliers, highlighting the mean (green star, indicated by the green arrow head), median, and density peak for a context length of 60 minutes and a prediction length of 90 minutes. See figures 13, 14 and 15 in E for the full analysis.
@@
Models 2h (metres)
TrAISformer 6945.2
MixTRAJ w/ finetuning (ours) 2528.4
MixTRAJ w/o finetuning (ours) 5407.6
Table 3: Comparison of trajectory prediction metrics between TrAISformer and Mixtral-TRAJ models. All measurements are in metres calculated with the Fréchet distance.
@@
The figure 6 highlights that the prediction error generally increases with the prediction distance, as shown by the comparison between (a) and (b). It is generally observed that the prediction errors are greater as the context length is reduced, and the prediction length is increased. On the other hand, we can see that, despite certain hallucinations which we will discuss later, the model seems to have generalised the nature of a trajectory, as shown in image (c) of figure 16 in appendix F.
@@
In contrast, as shown in Table 3, MixTRAJ demonstrates a substantial enhancement in predictive accuracy, outperforming the baseline TrAISformer model by $28.4\%$ without fine-tuning and by an impressive $174.7\%$ with fine-tuning on TrAISformer’s evaluation dataset. This result highlights that, even without fine-tuning on TrAISformer’s data and despite the limited representation of this specific region in our initial dataset, our baseline model performs better, suggesting it has successfully generalised the intrinsic nature of trajectory patterns. Notably, we opted to use the median error for model evaluation, whereas the initial TrAISformer implementation used the minimum error; we adapted this to ensure a more robust assessment.
@@ 5.3 Hallucination
Figure 9: An example of the phenomena we have called "hallucinations". A red line represents the context used to generate the green trajectory. A blue line represents the ground truth. See more example in the figures 17 and 18 in appendix G
One of the key reasons behind the occurrence of these hallucinations is linked to the nature of the Causal Language Model we used and its auto-regressive architecture. In an auto-regressive framework, each prediction is based solely on the previous tokens without any direct consideration of future points, making it prone to accumulating small errors over time. As these errors compound, especially in longer sequences, they can lead to unrealistic outputs such as abrupt turns or improbable position jumps.
@@ 6 Conclusion
Data Availability
Due to the sensitive nature of the data, which was collected using a proprietary system of the French Navy, we are unable to share the raw dataset. However, we have provided a detailed description of the data preparation procedure, ensuring that others can replicate the cleaning and tokenisation steps. If needed, we can also provide the code used for data cleaning. Moreover, the data can be recaptured dynamically, as we did during the testing phase, or exported from publicly available platforms such as MarineTraffic, allowing for similar experiments to be conducted with alternative datasets.
@@ Appendix F Multiple trajectories grid
Appendix G Hallucinations
Figure 17: More examples of hallucinations. The red point is the barycentre of the polygon formed by the green points at the ends of the green trajectories.
Figure 18: More examples of hallucinations.